(4)^1/2 + 2
It's asking me to say whether or not it is as real and rational, real and irrational, or complex. Ummm yeah how do I figure that out...she never went over it in class. :(

Answer by stanbon(48569)

4^(1/2)+2 = 2+2= 4
which is a rational number.
All rational numbers are real.
So I guess your teacher wants you to say it is real and rational.
Cheers,
Stan H.